# The Behrens–Fisher problem revisited

### v1

@@ 5.3 Welch versus compact Behrens–Fisher
The work of Nel et al. [28] is conditioned on the unknown variances $\sigma_{i}^{2}$. They express $T$ as a ratio $U/(V+W)$ with $U\sim\chi^{2}_{1}$ and $V,W$ independent Gamma variates, and evaluate the joint integral by a single change of variables, obtaining a $_{2}F_{1}$ kernel. We apply a MB identity to decouple $(gW_{1}+hW_{2})^{1/2}$, collapse the 2-fold integral to one contour; closing right gives the residue series, closing left yields the same hypergeometric pdf as [28]. Further, a separate $(S,U)$ polar transform leads to the Euler–Beta representation for the cdf. The MB approach also furnishes cumulant–generating functions and Ramanujan coefficients—absent in the work of [28]. From a computational standpoint, [28] called their density “computationally intractable.” We overcome the numeric hurdle by: (i) providing a terminating series when either $n_{i}$ is odd, (ii) showing $_{2}F_{1}$ in (29) is stable under the Möbius argument $\rho(t)$. Thus, we convert what [28] regarded as theoretical curiosity into a deployable routine. [28] do not discuss effective degrees of freedom. We relate the hypergeometric tail exponent to $\nu_{\text{tail}}=\nu_{1}+\nu_{2}$ and contrast it with Welch’s random $\nu_{W}$, clarifying why and where Welch inherits liberal or conservative bias. Our analytic heatmap quantify the maximum Type-I error inflation of Welch (up to +58 % for $n_{1}=n_{2}=10,\,r=4$) and locate the crossover surface $\nu_{W}=\nu_{\text{tail}}$. A summary of the comparison between Nel et al’s work and our work is summarized in Table 8. We also provide new analytic assets: (i) Ramanujan-series tails—usable for fast, high-accuracy upper tail $p$-values; (iii) Lugannani–Rice correction—back–checked against the exact tail, delivering $10^{-6}$ error for $|t|<0.8\min\{\nu_{1},\nu_{2}\}$; (iii) A cdf that permits likelihood maximization without numerical quadrature over $(W_{1},W_{2})$.
@@ 6 Concluding remarks
Given two independent samples $X_{1},\dots,X_{n_{1}}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}\mathcal{N}(\mu_{1},\sigma_{1}^{2})$, $Y_{1},\dots,Y_{n_{2}}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}\mathcal{N}(\mu_{2},\sigma_{2}^{2})$ with $\nu_{i}=n_{i}-1$ degrees of freedom, sample means $\bar{X}$, $\bar{Y}$, and sample variances $S_{1}^{2}$, $S_{2}^{2}$, $Z=\bar{X}-\bar{Y}$, $g=\sigma_{1}^{2}/n_{1}$, $h=\sigma_{2}^{2}/n_{2}$ and $r=\sigma_{1}^{2}/\sigma_{2}^{2}$, the Behrens–Fisher statistic $T=Z/\sqrt{S_{1}^{2}/n_{1}+S_{2}^{2}/n_{2}}$ to test $H_{0}\!:\mu_{1}=\mu_{2}$ has, under $H_{0}$ and $g\neq h$, a density that has resisted compact expression for several decades. We showed that a Mellin–Barnes decomposition of $(gW_{1}/\nu_{1}+hW_{2}/\nu_{2})^{1/2}$ collapses the joint integral over the independent $\chi^{2}$-variates $W_{i}=\nu_{i}S_{i}^{2}/\sigma_{i}^{2}$ to a single contour integral. Residue calculus yields a finite series whenever either $\nu_{i}$ is odd; an Euler–Beta identity provides a compact form for the density function $f_{T}(t)$in terms of Gauss’ hypergeometric function. Ramanujan’s master theorem aids exact tail analyses. Our result subsumes the results of Nel et al. [28] and extends it with a closed-form cdf and analytic tails. Critical values $(\alpha=0.10-0.001)$ computed over $2\leq n_{i}\leq 50$, $0.1\leq r\leq 10$ reveal the parameter surface on which Welch’s approximation [21] switches from conservative to liberal.

### v2

@@ 5.3 Welch’s approximation versus our unified approach
The work of Nel et al. (1990) is conditioned on the unknown variances $\sigma_{i}^{2}$. They express $T$ as a ratio $U/(V+W)$ with $U\sim\chi^{2}_{1}$ and $V,W$ independent Gamma variates, and evaluate the joint integral by a single change of variables, obtaining a $_{2}F_{1}$ kernel. We apply a MB identity to decouple $(gW_{1}+hW_{2})^{1/2}$, collapse the 2-fold integral to one contour; closing right gives the residue series, closing left yields the same hypergeometric pdf as Nel et al. (1990). Further, a separate $(S,U)$ polar transform leads to the Euler–Beta representation for the cdf. The MB approach also furnishes cumulant–generating functions and Ramanujan coefficients—absent in the work of Nel et al. (1990). From a computational standpoint, Nel et al. (1990) called their density “computationally intractable.” We overcome the numeric hurdle by: (i) providing a terminating series when either $n_{i}$ is odd, (ii) showing $_{2}F_{1}$ in (16) is stable under the Möbius argument $\rho(t)$. Thus, we convert what Nel et al. (1990) regarded as theoretical curiosity into a deployable routine. Nel et al. (1990) do not discuss effective degrees of freedom. We relate the hypergeometric tail exponent to $\nu_{\text{tail}}=\nu_{1}+\nu_{2}$ and contrast it with Welch’s random $\nu_{W}$, clarifying why and where Welch inherits liberal or conservative bias. Our analytic heatmap quantify the maximum Type-I error inflation of Welch (up to +58 % for $n_{1}=n_{2}=10,\,r=4$) and locate the crossover surface $\nu_{W}=\nu_{\text{tail}}$. A comparison between the work of Nel et al. (1990) and our work is summarized in the following paragraph.
Nel et al. (1990) derived the univariate Behrens–Fisher pdf in a single Gauss hypergeometric expression; in contrast, our work develops two algebraically equivalent forms—namely, a terminating or convergent residue series and a hypergeometric representation. Nel et al. (1990) did not obtain a closed-form cdf, whereas we provide a one-dimensional Euler–Beta integral that reduces to the Appell $F_{1}$ function and yields an exact cdf valid for all $(n_{1},n_{2})$. Their work offers neither tail expansions nor analytic error bounds, while we present Ramanujan-based coefficients of order $t^{-(2m+1)}$ along with a computable remainder. They also did not develop saddle-point diagnostics; our work includes a Lugannani–Rice approximation based on explicit cumulants from the Mellin–Barnes generator, achieving uniform $O(t^{-4})$ accuracy away from boundary regions. Nel et al. (1990) provided no critical-value tables, whereas we supply exact 0.10–0.001 quantiles for arbitrary $(n_{i},r)$. Their article did not benchmark against Welch’s test; we include a heatmap and crossover contour identifying the sign-change region of Welch’s statistic relative to the exact solution. Finally, although Nel et al. (1990) sketched a multivariate extension via Hotelling’s $T^{2}$, we defer the multivariate development to future work while noting that our univariate derivations generalize naturally. We also provide new analytic assets: (i) Ramanujan-series tails—usable for fast, high-accuracy upper tail $p$-values; (iii) Lugannani–Rice correction—back–checked against the exact tail, delivering $10^{-6}$ error for $|t|<0.8\min\{\nu_{1},\nu_{2}\}$; (iii) A cdf that permits likelihood maximization without numerical quadrature over $(W_{1},W_{2})$.
@@ 6 Concluding remarks
Given two independent samples $X_{1},\dots,X_{n_{1}}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}\mathcal{N}(\mu_{1},\sigma_{1}^{2})$, $Y_{1},\dots,Y_{n_{2}}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}\mathcal{N}(\mu_{2},\sigma_{2}^{2})$ with $\nu_{i}$ degrees of freedom, sample means $\bar{X}$, $\bar{Y}$, and sample variances $S_{1}^{2}$, $S_{2}^{2}$, $Z=\bar{X}-\bar{Y}$, $g=\sigma_{1}^{2}/n_{1}$, $h=\sigma_{2}^{2}/n_{2}$ and $r=\sigma_{1}^{2}/\sigma_{2}^{2}$, the Behrens–Fisher statistic $T=Z/\sqrt{S_{1}^{2}/n_{1}+S_{2}^{2}/n_{2}}$ to test $H_{0}\!:\mu_{1}=\mu_{2}$ has, under $H_{0}$ and $(\sigma_{1}^{2},\sigma_{2}^{2})$ unknown, $\sigma_{1}^{2}\neq\sigma_{2}^{2}$, a density that has resisted compact expression for several decades. We showed that a Mellin–Barnes decomposition of $(gW_{1}/\nu_{1}+hW_{2}/\nu_{2})^{1/2}$ collapses the joint integral over the independent $\chi^{2}$-variates $W_{i}=\nu_{i}S_{i}^{2}/\sigma_{i}^{2}$ to a single contour integral. Residue calculus yields a finite series whenever either $\nu_{i}$ is odd; an Euler–Beta identity provides a compact form for $f_{T}(t)$in terms of Gauss’ hypergeometric function. Ramanujan’s master theorem aids exact tail analyses. This unified approach thereby makes residue calculus, Euler–Beta reduction, and Ramanujan-type asymptotics all available within a single framework. Our result subsumes the pdf of Nel et al. (1990) and extends it with a closed-form cdf and analytic tails. Critical values $(\alpha=0.10-0.001)$ computed over $2\leq n_{i}\leq 50$, $0.1\leq r\leq 10$ reveal the parameter surface on which Welch (1947) approximation switches from conservative to liberal.
Disclosure of interest
The authors report there are no competing interests to declare.